40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 1413     Entry time: Tue May 10 11:53:48 2011
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: no real progress 

In a heightened state of despair-induced laziness, I forgot to elog what I did yesterday:

Last week I discovered that there was something wrong with the primary PDH box, so I started there. I was unable to take a transfer function of the overall servo because the output kept railing (I tried adjusting the DC offset). I then noticed that the second switch I had added to optionally remove the DC-gain-limiting resistor in the first, initially non-switchable TF stage (U7) had come loose. I had put this in so that we could get 1/f2 at low frequency, since the baseline design of the servo only allows 1/f. I took a transfer function of this stage alone, and it looked awful. I removed the extra switch and put back the 1k SMT resistor to limit the DC gain, and this made the TF look like it should (a pole and a zero). I am guessing that the OP27 just couldn't handle what we were trying to do with it.

I reinstalled the servo and the cavity wouldn't lock. This time, though, I was able to get it to lock by simply adding flat gain with an SR560 between the demod electronics and the servo input. I concluded that we just didn't have enough gain now that the optical power has been lowered by >20 dB. I needed G = 500 on the SR560, so I changed R23 (between the two TF stages U7 & U4) from 110k to 220, adding roughly the right amount. Note that this is weird, since we didn't lower the optical power by this much... something is fishy.

I measured the OLTF of the primary loop to ensure that we were back at roughly the standard operating configuration. I found that the UGF was at ~10 kHz, which is just about right (the servo gain could probably be increased slightly without oscillation).


With everything running again, I locked the second direction so that I could measure the gyro noise. The OL gain of the secondary loop is now much lower than it was---since I haven't added servo gain---but the UGF is >100 Hz, so we can trust the LF spectrum we get. 

The gyro noise now exhibits a strange sort of f-1/2 noise at lower frequencies, and is considerably higher (~10-100x) than it was before we moved the FIs. I suspect that this noise is from a different source than the old LF "hump", and hopefully that contribution is now lower. The task is to figure out what is causing the new noise. I am suspicious of the amount of gain I had to add to the primary servo, so I think it's wise to go back and check our modulation depth, etc. to ensure something's not screwed up.


ELOG V3.1.3-